OpenStudy (anonymous):

Just need someone to check my answers to make sure I'm on the right track y=x^2+8x-14 is a.... Parabola. Form y=(x+4)^2 + 30 Vertex (-4,-30) x^2=12+4y^2-8y is a... Hyperbola. Form x^2/8 - (y-1)^2 /2 =1 Center (0,1)

OpenStudy (anonymous):

I know I identified the problems correctly as Parabola, Ellipse, Circle, etc... just the part that said "Form" and had a blank confused me

OpenStudy (tkhunny):

#1 (x+4)^2 + 30 = x^2 + 8x + 16 + 30 = x^2 + 8x + 46 Seems like something went wrong, there. Try Completing the Square, again.

OpenStudy (anonymous):

for instance, \[4x ^{2} -16x = -9y ^{2} -18y + 11\] is an Ellipse, right? How am I supposed to get this into Ellipse form

OpenStudy (anonymous):

Checking #1 right now... I checked it on Mathway. Not sure how that happened

OpenStudy (tkhunny):

Looks like #2 is good! Good work.

OpenStudy (anonymous):

I checked #1 on Mathway, clicked "Find Vertex form" and it came up y=(x+4)^2 - 30. Put PLUS instead of MINUS 30. Does that make it correct?

OpenStudy (tkhunny):

You tell me. We had -14. We added 16. We should correct with another -16. -14 - 16 = ??

OpenStudy (anonymous):

30! Ok, that makes sense! Could you help me on that 3rd one I wrote above? The Ellipse

OpenStudy (anonymous):

Nevermind, I think I got it! It's \[\frac{ (x-2)^{2} }{ 9 }+\frac{ (y-1)^{2} }{ 4}=1\]